Assume a $40,000 investment and the following cash flows for two alternatives. Year Investnent x Investment Y 9 6,000 8,000 9,000 17,000 20,000 $ 15,000 20,000 10,000 late the payback for investment X and Y. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Payback nvestment X years nvestment Y years b. Which alternative would you select under the payback method? Investment Х. Investment Y
Answer #1

a)

Investment X:

Cumulative cash flow for year 0 = -40,000

Cumulative cash flow for year 1 = -40,000 + 6,000 = -34,000

Cumulative cash flow for year 2 = -34,000 + 8,000 = -26,000

Cumulative cash flow for year 3 = -26,000 + 9,000 = -17,000

Cumulative cash flow for year 4 = -17,000 + 17,000 = 0

Payback period = 4 years

Investment Y:

Cumulative cash flow for year 0 = -40,000

Cumulative cash flow for year 1 = -40,000 + 15,000 = -25,000

Cumulative cash flow for year 2 = -25,000 + 20,000 = -5,000

Cumulative cash flow for year 3 = -5,000 + 10,000 = 5,000

5,000 / 10,000 = 0.5

Payback period = 2.5 years

2)

Investment Y as it has a lower payback period.